A company just paid a dividend of 2.30 to its shareholders. It estimates that future growth will be at 2%. What is the value of the stock, if you are looking for an 8% return on your investment?

What led you to this answer?
Answer #1

Div1 = $2.30 * (1 + 2%) = $2.346

P_0 = \frac{2.346}{0.08 - 0.02}

P_0 = \frac{2.346}{0.06}

P_0 = 39.10